OpenStudy (anonymous):

Integrate the indefinite integral: \[\int\limits_{}^{}\frac{ \sin ^{2}x+\cos ^{2}x }{ cscx} dx\]

ganeshie8 (ganeshie8):

knw ur trig :)

ganeshie8 (ganeshie8):

\(\large \sin^2 x + \cos^2 x = 1\)

ganeshie8 (ganeshie8):

\(\large \int\limits_{}^{}\frac{ \sin ^{2}x+\cos ^{2}x }{ cscx} dx = \int \frac{1}{\csc x} dx = \int \sin x dx = ? \)